The market for USB flash drives in Country C is perfectly competitive and is in equilibrium.
Domestic demand is given by Qd = 300 – 4P and domestic supply is given by Qs = 2P.
The world price for flash drives is $20.

The government of country C imposes a tariff of $20 on all imported flash drives.

Before the tariff was imposed, country C imported MFree flash drives.
After the tariff is imposed, country C now imports MTrf flash drives.


What is MFree and MTrf?

Question 1 options:

1)

MFree = 220 units, MTrf = 60 units

2)

MFree = 180 units, MTrf = 140 units

3)

MFree = 140 units, MTrf = 60 units

4)

MFree = 180 units, MTrf = 60 units

5)

MFree = 220 units, MTrf = 140 units
Answer #1

supply (Qs = 20) E inchal Equubrien 50 760- Od = 300-up Qd = 3 Putt = 40 -------- PW = $20 PO 1400 At Pw = $20, the Quantity